# Modeling the dynamics of an oil drop driven by a surface acoustic wave in the underlying substrate

### v3

@@ 4 Theoretical results
In this section, we discuss the predictions of the model developed in Sec. 3, and compare them with the experimental results. In Sec. 4.1, we present results for a few chosen reference cases, and then we discuss the influence of model parameters, in particular the acoustic amplitude, $A$, on the results. In Sec. 4.2 we discuss (approximate) traveling-wave solutions.
@@ 4.1 Main features of the results
In our simulations, we solve Eq. (29) numerically using COMSOL™; see Appendix C for the description of the implementation. As the initial condition, we consider the two–dimensional parabolic drop of cross-sectional area
@@
which remains constant during the drop evolution. The drop height parameter $\tilde{h}_{d}$ is obtained from the known experimental (3D) drop volume, ${\cal V}_{d}=\ell^{3}\tilde{\cal V}_{d}$, as $\tilde{h}_{d}=2\tilde{\cal V}_{d}/(\pi\tilde{r}_{d}^{2})$, using the formula for the volume $\tilde{\cal V}_{d}$ of a 3D parabolic cap of base radius $\tilde{r}_{d}$ and height $\tilde{h}_{d}$. The 2D parabolic initial condition is centered at $\tilde{x}=\tilde{x}_{d}$, and is written as
$\tilde{h}(\tilde{x},0)=\left\{\begin{array}[]{ll}\tilde{h}_{p},&\tilde{x}<\tilde{x}_{\rm r}(0),\\ \left(\tilde{h}_{d}-\tilde{h}_{p}\right)\left[1-\left(\frac{\tilde{x}-\tilde{x}_{d}}{\tilde{r}_{d}}\right)^{2}\right]+\tilde{h}_{p},&\tilde{x}_{\rm r}(0)\leq\tilde{x}\leq\tilde{x}_{\rm f}(0),\\ \tilde{h}_{p},&\tilde{x}>\tilde{x}_{\rm f}(0),\end{array}\right.$ (33)
where $\tilde{x}_{\rm r}(0)=\tilde{x}_{d}-\tilde{r}_{d}$ and $\tilde{x}_{\rm f}(0)=\tilde{x}_{d}+\tilde{r}_{d}$ are the initial positions of the rear and front contact lines of the drop, respectively (see Fig. 4). Here, $\tilde{h}_{p}$ ($\ll\tilde{h}_{d}$) is the thickness of a precursor film, introduced to avoid the stress singularity associated with a moving contact line. It is well known (see, e.g. Diez & Kondic (2002)), that the step size $\Delta\tilde{x}$ in the spatial discretization should be similar to $\tilde{h}_{p}$ to ensure numerical convergence, and therefore we use $\Delta\tilde{x}=\tilde{h}_{p}$ in our simulations. The specific value assigned to $\tilde{h}_{p}$ has only a minor influence on the results; e.g., if $\tilde{h}_{p}$ is doubled or halved, the change of spreading speed or drop height is just 1-2%. All reported results are obtained using $\tilde{h}_{p}=0.01$.
$\psi(\tilde{x},\tilde{h})=\left\{\begin{array}[]{ll}1,&\tilde{x}<\tilde{x}_{1}^{\ast}(\tilde{t})\,,\\ e^{-2\left[\tilde{k}_{s,i}(\tilde{x}-\tilde{x}_{1}^{\ast}(\tilde{t}))+\tilde{K}_{z}\left(\tilde{h}-\tilde{h}^{\ast}\right)\right]},&\tilde{x}_{1}^{\ast}(\tilde{t})\leq\tilde{x}\leq\tilde{x}_{2}^{\ast}(\tilde{t})\,,\\ e^{-2\tilde{k}_{s,i}\left[\tilde{x}_{2}^{\ast}(\tilde{t})-\tilde{x}_{1}^{\ast}(\tilde{t})\right]},&\tilde{x}>\tilde{x}_{2}^{\ast}(\tilde{t})\,,\end{array}\right.$ (34)
where $\tilde{x}_{1}^{\ast}$, $\tilde{x}_{2}^{\ast}$ are defined by $\tilde{h}(\tilde{x}_{1}^{\ast},\tilde{t})=\tilde{h}(\tilde{x}_{2}^{\ast},\tilde{t})=\tilde{h}^{\ast}$, and the SAW is assumed to propagate from $\tilde{x}<\tilde{x}_{1}^{*}(\tilde{t})$ in the direction of increasing $\tilde{x}$. In line with our discussion in Sec. 3, in simulations we choose the cutoff film thickness $\tilde{h}^{*}$ based on the thickness below which the bulk acoustic streaming becomes weak and mass transport is governed by other mechanisms (e.g., Rayleigh streaming and acoustic pressure (Rezk et al., 2012, 2014)), $\tilde{h}^{\ast}\approx(\lambda_{\rm oil}/4)/\ell=17\mu{\rm m}/\ell$.
We commence our analysis by considering the experimental case where an oil drop of volume $8$ mm3 is under the excitation of the $20$ MHz SAW.
@@
$\mathcal{S}$ $(\rho_{0}\ell A^{2}\omega^{2})/\gamma$ $0.788-72.88$ Ratio of acoustic force and surface tension
@@
$\tilde{\mathcal{C}}$ $\tilde{K}_{z}\tilde{C}_{x}-\tilde{k}_{\rm s,i}\tilde{C}_{z}$ $-0.11$ Effect of non-conservative SAW force
@@
$\tilde{\mathcal{A}}_{d}$ $(4/3)\tilde{h}_{d}\tilde{r}_{d}$ $2.12$ Area of drop
$\tilde{h}_{d}$ $(2\tilde{\mathcal{V}}_{d})/(\pi\tilde{r}_{d}^{2})$ $0.497$ Height of drop
$\tilde{r}_{d}$ $3.2$ Radius of drop
$\tilde{h}_{p}$ $0.01$ Precursor film thickness
Table 2: Dimensionless parameters used in the simulation of the experiment for a drop volume of $\mathcal{V}_{d}=8$ mm3 driven by a SAW with $A$ in the range from Table 1, which determines the range of $\mathcal{S}$.
Figure 5 shows the evolution of the thickness profile obtained by solving Eq. (29) for $A=1.3,2,6,10$ nm subject to the initial condition given by Eq. (33), and boundary conditions fixing the film thickness to $h_{p}$ and requiring vanishing derivatives at the domain boundaries; the other parameter values are shown in Table 2. For the low-amplitude cases (e.g. $A<2$  nm, Fig. 5(a) and (b)), we see that capillary and gravitational forces dominate, leading to a continual spreading of the drop. The acoustic force is able to slightly drive the drop along the direction of propagation of the SAW, as is apparent by the movement of the two contact lines. In contrast, in the high-amplitude cases (e.g., $A>2$ nm, Fig. 5(c) and (d)), the acoustic force plays a dominant role in governing the dynamics. We see that the evolution of the drop can be described by two regimes: an initial transient stage where the drop takes on a new shape, followed by a period of translation across the domain at near-constant speed, consistent with experimental results (see Fig. 3); as $A$ increases, the duration of the transient stage decreases. The fluid region near the rear contact line, where the pressure gradient is largest, serves as a “snow plow” that pushes the bulk drop along the direction of propagation of the SAW. The result is a near traveling wave solution at long times, discussed further in Sec. 4.2. The subtle, but continuous, decrease of the maximum height of the drop during this period is due to a thin trailing film (of approximate thickness $h^{\ast}$) that is left behind the bulk drop. We also note formation of a trailing foot forming spontaneously behind the main body of the drop for amplitudes $A>2$ nm. This feature travels together with the main body of the droplet, in contrast to the trailing film, which stays behind.
(a) $A=1.3$ nm
@@
Figure 5: Evolution of thickness profile for different values of $A$. The black solid lines correspond to the initial condition, the dashed black lines to the profiles every time step $\Delta t=1$ s, and the solid blue line to $t=6~{}$s. The red line stands for the effective dimensionless pressure $\tilde{\mathcal{P}}$ at $t=6$ s. Note that the dip in the pressure curve corresponds to the dramatic change of curvature of the fluid interface at the front contact line.
@@
Focusing next on the spreading speed, Fig. 6 plots the relative speed of spreading for both experiments and simulations. The numerical results (solid lines) show a monotonic increase in spreading speed as $A$ increases, consistent with the experimental results (symbols). Note the distinct regimes in which the drops evolve, evident from the apparent change of slope of the curves, especially prominent in large-amplitude cases. Figure 6(b) shows the numerically-calculated drop width, ${\rm w}=x_{2}^{\ast}-x_{1}^{\ast}$, as a function of time for different values of $A$. We once again emphasize the different behaviors exhibited by low- and high-amplitude solutions. For low-amplitude cases, the front contact line of the drop moves faster than the rear contact line, causing the drop to become thinner and wider. Conversely, for the high-amplitude cases, $\rm{w}$ initially increases and then remains nearly constant, or even decreases slowly.
@@
Figure 6(a) illustrates a quantitative difference between experiment and simulations: assuming that $A=A_{n}$ (see also the relevant discussion in Sec. 3), the simulated drops travel at a slower pace. While it is not clear at this point where this difference is coming from, some possible reasons, in addition to possible differences between $A$ and $A_{n}$ (Royer & Dieulesaint, 1996), include: (i) the possible relevance of inertial effects that are not included in the model at order $M_{a}$ (the solution $\mathbf{v}_{1}$(Zarembo, 1971; Orosco & Friend, 2021; Dubrovski et al., 2023); (ii) the omission of contributions from the boundary layer flow near the solid surface, i.e., the Rayleigh law of streaming; (iii) the capillary waves that appear at the free surface of the film (Lighthill, 1978; Morozov & Manor, 2017); (iv) the use of a two-dimensional model to understand a three-dimensional system; and (v) the neglect of the ultrasonic wave reflection off the free surface of the liquid. We have verified that the numerical results are independent of the precise values assigned to $h^{*}$ and $h_{p}$, suggesting that neither boundary layer thickness nor precursor thickness are relevant here. We have also carried out some preliminary simulations of a three-dimensional oil drop, which suggest that the two-dimensional model used in the present work is not responsible for the quantitative differences between experiments and simulations (we leave the discussion of three-dimensional effects for future work). Clearly, more work will be needed to uncover the sources of the quantitative differences between experiments and simulations.
@@
Figure 6: (a) Comparison between numerical (solid lines, amplitude values $A$ as labeled, color-coded) and experimental results (solid dots, for measured amplitude values $A_{n}=1.69\,$nm (blue), 1.301.301.30\,1.30nm (red), 1.041.041.04\,1.04nm ( olive)). (b) Numerically-calculated drop width data plotted as a function of time for several $A$-values (color-matched with panel (a) where appropriate).
@@
Figure 7: Long time (calculated at $t=9$ s) values of the front speed, $\rm{v}_{\rm f}$ (in mm/s, blue filled circles), maximum thickness, $h_{\rm Max}$ (mm, red filled squares), and drop width, $\rm w$ (cm, black filled circles), as functions of $A$. Note that $\rm v_{\rm f}$ is plotted using the left $y$-axis while $\rm w$ and $\rm h_{\rm{Max}}$ are plotted using the right $y$-axis. The points correspond to the raw data, and the solid lines connecting them guide the eye.
@@
To summarize our numerical results, Fig. 7 plots the numerically-predicted values of the front velocity, ${\rm v}_{\rm f}$, the maximum drop height, $h_{\rm Max}$, and the drop width, $\rm{w}$, all calculated at $t=9$ s, as a function of SAW amplitude $A$. This figure illustrates the transition between the two types of dynamics observed: for low and moderate $A$-values ($A\leq 2$ nm) we observe near-symmetric drop spreading due primarily to capillary forces, however, for large values ($A>2$ nm) the front speed increases approximately quadratically as $A$ increases, while the maximum height/width of the drop increases/decreases approximately linearly with respect to the amplitude value.
@@ 4.2 Traveling wave solution
where (constant) $\tilde{\rm{w}}$ is the width of the moving drop. Here, we will consider that the rear (front) contact lines of the traveling drop are given by $\tilde{x}_{1}^{\ast}$ ($\tilde{x}_{2}^{\ast}$) (see Fig. 4), since the SAW force that drives it is felt only where $\tilde{h}>\tilde{h}^{\ast}$. Then, we define $\tilde{\xi}=\tilde{x}-\tilde{x}_{1}^{\ast}(\tilde{t})=\tilde{x}-\tilde{U}\tilde{t}$ and assume that $\tilde{h}(\tilde{x},\tilde{t})=\tilde{h}(\tilde{\xi})$, so that Eq. (29) becomes (see also Buckingham et al. (2003); Perazzo & Gratton (2004); Giacomelli et al. (2016) for similar studies without SAW forces)
@@
$\psi(\tilde{\xi},\tilde{h})=e^{-2[\tilde{k}_{\rm s,i}\tilde{\xi}+\tilde{K}_{z}(\tilde{h}-\tilde{h}^{\ast})]}.$ (37)
@@
[…]
@@
We developed an iterative scheme to perform the numerical integration of Eq. (38), under the conditions given by Eq. (40). We start by guessing values of $\tilde{h}^{\prime\prime}(0)$, $\tilde{U}$ and $\tilde{\xi}_{\rm f}$ to perform the integration of Eq. (38) and modify them accordingly until these conditions are satisfied within a small relative error (typically, $10^{-7}$ or even smaller). After obtaining convergence for a given $\tilde{A}$, we use these converged values of $\tilde{h}^{\prime\prime}(0)$, $\tilde{U}$ and $\tilde{\xi}_{\rm f}$ as new guess values for $\tilde{A}+\Delta\tilde{A}$ (continuation; $\Delta\tilde{A}$ has to be very small to reach convergence).
@@
This approximate formulation shows that the simple traveling-wave solution can capture, at least qualitatively, certain features of the experimental and numerical drops. Clearly, this description cannot account for the rear region of the drop (see Figs. 5 and 8(a)) because the values of $h^{\prime}$ and $h^{\prime\prime\prime}$ at the left contact line cannot be determined a priori, and both have been set to zero (for lack of better choice) in the traveling wave calculations.
Figure 8: (a) Thickness profile of the traveling wave solution for: $A=4.5$ nm, $6$ nm, $8$ nm, $10$ nm (b) Front speed, ${\rm v}_{\rm f}$ (mm/s, blue line), drop width, $\rm w$ (mm, black line), and maximum height, $h_{\rm max}$ (mm, red line), of the traveling drop as a function of $A$; front speed is plotted on the left $y$-axis while drop width and maximum height are on the right $y$-axis.

### v4

@@ 4 Theoretical results
In this section, we discuss the predictions of the model developed in Sec. 3, and compare them with the experimental results. In Sec. 4.1, we present results for a few chosen reference cases, and then we discuss the influence of model parameters, including the acoustic amplitude, $A$, the droplet volume $\mathcal{V}_{\rm d}$ and the droplet kinematic viscosity $\nu$, on the results. In Sec. 4.2 we discuss (approximate) traveling-wave solutions.
@@ 4.1 Main features of the results
In our simulations, we solve Eq. (29) numerically using COMSOL™; see Appendix C for the description of the implementation. As the initial condition, we consider the two–dimensional parabolic drop of (dimensionless) cross-sectional area
$\tilde{{\cal A}}_{\rm d}=2\int_{0}^{\tilde{r}_{\rm d}}\tilde{h}_{\rm d}\left[1-\left(\frac{\tilde{x}}{\tilde{r}_{\rm d}}\right)^{2}\right]\,d\tilde{x}=\frac{4}{3}\tilde{h}_{\rm d}\tilde{r}_{\rm d},$ (33)
(recall that tilded quantities are dimensionless), which remains constant during the drop evolution. The drop height parameter $\tilde{h}_{\rm d}$ is obtained from the known experimental (3D) drop volume, ${\cal V}_{\rm d}=\ell^{3}\tilde{\cal V}_{\rm d}$, as $\tilde{h}_{\rm d}=2\tilde{\cal V}_{\rm d}/(\pi\tilde{r}_{\rm d}^{2})$, using the formula for the volume $\tilde{\cal V}_{\rm d}$ of a 3D parabolic cap of base radius $\tilde{r}_{\rm d}$ and height $\tilde{h}_{\rm d}$. The 2D parabolic initial condition is centered at $\tilde{x}=\tilde{x}_{\rm d}$, and is written as
$\tilde{h}(\tilde{x},0)=\left\{\begin{array}[]{ll}\tilde{h}_{\rm p},&\tilde{x}<\tilde{x}_{\rm r}(0),\\ \left(\tilde{h}_{\rm d}-\tilde{h}_{\rm p}\right)\left[1-\left(\frac{\tilde{x}-\tilde{x}_{{\rm d}}}{\tilde{r}_{\rm d}}\right)^{2}\right]+\tilde{h}_{\rm p},&\tilde{x}_{\rm r}(0)\leq\tilde{x}\leq\tilde{x}_{\rm f}(0),\\ \tilde{h}_{\rm p},&\tilde{x}>\tilde{x}_{\rm f}(0),\end{array}\right$. (34)
where $\tilde{x}_{\rm r}(0)=\tilde{x}_{\rm d}-\tilde{r}_{\rm d}$ and $\tilde{x}_{\rm f}(0)=\tilde{x}_{\rm d}+\tilde{r}_{\rm d}$ are the initial positions of the rear and front contact lines of the drop, respectively (see Fig. 5). Here, $\tilde{h}_{\rm p}$ ($\ll\tilde{h}_{\rm d}$) is the thickness of a precursor film, introduced to avoid the stress singularity associated with a moving contact line. It is well known (see, e.g. Diez & Kondic (2002)), that the step size $\Delta\tilde{x}$ in the spatial discretization should be similar to $\tilde{h}_{\rm p}$ to ensure numerical convergence, and therefore we use $\Delta\tilde{x}=\tilde{h}_{\rm p}$ in our simulations. The specific value assigned to $\tilde{h}_{\rm p}$ has only a minor influence on the results; e.g., if $\tilde{h}_{\rm p}$ is doubled or halved, the change of spreading speed or drop height is just 1-2%. All reported results are obtained using $\tilde{h}_{\rm p}=0.01$.
@@
[…]
@@
$\mathcal{S}$ $(\rho_{0}\ell A^{2}\omega^{2})/\gamma$ 0.78872.880.788-72.88 Ratio of acoustic force and surface tension
@@
$\tilde{\mathcal{C}}$ $\tilde{K}_{z}\tilde{C}_{x}-\tilde{k}_{\rm s,{\rm i}}\tilde{C}_{z}$ 0.11-0.11 Effect of non-conservative SAW force
@@
$\tilde{\mathcal{A}}_{\rm d}$ $(4/3)\tilde{h}_{\rm d}\tilde{r}_{\rm d}$ $2.12$ Area of drop
$\tilde{h}_{\rm d}$ $(2\tilde{\mathcal{V}}_{\rm d})/(\pi\tilde{r}_{\rm d}^{2})$ $0.497$ Height of drop
$\tilde{r}_{\rm d}$ $3.2$ Radius of drop
$\tilde{h}_{\rm p}$ $0.01$ Precursor film thickness
[…]
Figure 6 shows the evolution of the thickness profile obtained by solving Eq. (29) for $A=1.44,2,6,10$ nm subject to the initial condition given by Eq. (34), and boundary conditions fixing the film thickness to $h_{\rm p}$ and requiring vanishing derivatives at the domain boundaries; the other parameter values are shown in Table 2. For the low-amplitude cases (e.g. $A<2$  nm, Fig. 6(a) and (b)), we see that capillary and gravitational forces dominate, leading to a continual spreading of the drop. The acoustic force is able to slightly drive the drop along the direction of propagation of the SAW, as is apparent by the movement of the two contact lines. In contrast, in the high-amplitude cases (e.g., $A>2$ nm, Fig. 6(c) and (d)), the acoustic force plays a dominant role in governing the dynamics. We see that the evolution of the drop can be described by two regimes: an initial transient stage where the drop takes on a new shape, followed by a period of translation across the domain at near-constant speed, consistent with experimental results (see Fig. 3); as $A$ increases, the duration of the transient stage decreases. The fluid region near the rear contact line, where the pressure gradient is largest, serves as a “snow plow” that pushes the bulk drop along the direction of propagation of the SAW. The result is a near traveling wave solution at long times, discussed further in Sec. 4.2. The subtle, but continuous, decrease of the maximum height of the drop during this period is due to a thin trailing film (of approximate thickness $h^{\ast}$) that is left behind the bulk drop. We also note formation of a trailing foot forming spontaneously behind the main body of the drop for amplitudes $A>2$ nm. This feature travels together with the main body of the droplet, in contrast to the trailing film, which stays behind. We direct the interested reader to the supplementary materials where animations for each of these four cases are shown (movie 4, movie 5, movie 6, movie 7).
(a) $A=1.44$ nm
@@
Figure 6: Evolution of thickness profile for different values of $A$. The black solid lines correspond to the initial condition, the dashed black lines to the profiles every time step $\Delta t=1$ s, and the solid blue line to t=6t=6~s. The red line stands for the effective dimensionless pressure $\tilde{\mathcal{P}}$ at $t=6$ s. Note that the dip in the pressure curve corresponds to the dramatic change of curvature of the fluid interface at the front contact line.
@@
Focusing next on the spreading speed, Fig. 7 plots the relative speed of spreading for both experiments and simulations. The numerical results (solid lines) show a monotonic increase in spreading speed as $A$ increases, consistent with the experimental results (symbols). Note the distinct regimes in which the drops evolve, evident from the apparent change of slope of the curves, especially prominent in large-amplitude cases. Figure 7(b) shows the numerically-calculated drop width, ${\rm w}=x_{2}^{\ast}-x_{1}^{\ast}$, as a function of time, for different values of $A$. We once again emphasize the different behaviors exhibited by low- and high-amplitude solutions. For low-amplitude cases, the front contact line of the drop moves faster than the rear contact line, causing the drop to become thinner and wider. Conversely, for the high-amplitude cases, $\rm{w}$ initially increases and then remains nearly constant, or even decreases slowly.
@@
Figure 7(a) illustrates a quantitative difference between experiment and simulations: assuming that $A=A_{\rm n}$ (see also the relevant discussion in Sec. 3), the simulated drops travel at a slower pace. While it is not yet clear why this difference arises, some plausible reasons, in addition to possible differences between $A$ and $A_{\rm n}$ (Royer & Dieulesaint, 1996), include: (i) the potential relevance of inertial effects that are not included in the model at order $M_{a}$ (the solution $\mathbf{v}_{1}$(Zarembo, 1971; Orosco & Friend, 2021; Dubrovski et al., 2023); (ii) the omission of contributions from the boundary layer flow near the solid surface, i.e., the Rayleigh law of streaming; (iii) the capillary waves that appear at the free surface of the film (Lighthill, 1978; Morozov & Manor, 2017); (iv) the use of a two-dimensional model to understand a three-dimensional system; and (v) the neglect of the ultrasonic wave reflection off the free surface of the liquid. We have verified that the numerical results are independent of the precise values assigned to $h^{*}$ and $h_{\rm p}$, suggesting that neither boundary layer thickness nor precursor thickness are relevant here. We have also carried out some preliminary simulations of a three-dimensional oil drop, which suggest that the two-dimensional model used in the present work is not responsible for the quantitative differences between experiments and simulations (we leave the discussion of three-dimensional effects for future work). Clearly, more work will be needed to uncover the sources of the quantitative differences between experiments and simulations.
@@
Figure 7: (a) Comparison between numerical (solid lines, amplitude values AA as labeled, color-coded) and experimental results (solid dots, for measured amplitude values $A_{\rm n}=1.60\,$nm (blue), 1.441.44\,nm (red), 0.950.95\,nm ( olive)). (b) Numerically-calculated drop width data plotted as a function of time for several AA-values (color-matched with panel (a) where appropriate).
@@
Figure 8: Long time (calculated at $t=9$ s) values of the front speed, $\rm{v}_{\rm f}$ (in mm/s, blue filled circles), maximum thickness, $h_{\rm Max}$ ($m$, red filled squares), and drop width, $\rm w$ (cm, black filled circles), as functions of $A$. Note that $\rm v_{\rm f}$ is plotted using the left $y$-axis while $\rm w$ and $h_{\rm{Max}}$ are plotted using the right $y$-axis. The points correspond to the raw data, and the solid lines connecting them guide the eye.
@@
To summarize the numerical results presented so far, Fig. 8 plots the numerically-predicted values of the front speed, ${\rm v}_{\rm f}$, the maximum drop height, $h_{\rm Max}$, and the drop width, $\rm{w}$, all calculated at $t=9$ s, as a function of SAW amplitude $A$. This figure illustrates the transition between the two types of dynamics observed: for low and moderate $A$-values ($A\leq 2$ nm) we observe near-symmetric drop spreading due primarily to capillary forces, however, for large values ($A>2$ nm) the front speed increases approximately quadratically as $A$ increases, while the maximum height/width of the drop increases/decreases approximately linearly with respect to the amplitude value.
Before concluding this section, we briefly discuss how the results depend on the oil viscosity. We have also carried out simulations using different drop volumes; while the results are consistent with the experimental ones, shown in Fig. 4b, the influence of the change of fluid volume is, however, weaker than in the experiments (the results not shown for brevity). When varying the kinematic viscosity, $\nu$ (assuming $\nu^{\prime}=\nu$), without acoustic forcing ($A=0$), viscosity only (linearly) affects the time scale. With acoustics, however, the dynamics change in a nonlinear fashion, see Fig. 9. The mechanism behind this nonlinear dependence is traced to the prefactor of the nonconservative term in Eq. (29), which decreases as $\tilde{K}_{z}$ (defined in Eq. (47)) grows and $\tilde{\mathcal{C}}$ (defined in Eq. (23)) decreases in magnitude, yielding a relatively large attenuation length. This substantially weakens the forcing at the front, while nearly balancing gravitational and capillary forces at the rear, effectively stalling the drop motion for large viscosity values, in qualitative agreement with the experimental findings of Fig. 4(a).
(a) $\nu=\nu^{\prime}=50\,\text{cSt}$
(b) $\nu=\nu^{\prime}=100\,\text{cSt}$
(c) $\nu=\nu^{\prime}=500\,\text{cSt}$
Figure 9: Evolution of thickness profile for $A=2$ nm for drops of different viscosities, $\nu=\nu^{\prime}=50,100,500$ cSt, indicated in each figure part. The black solid lines correspond to the initial condition, the dashed black lines to the profiles every time step $\Delta t=2$ s, and the solid blue line to t=10t=10~s. The red line denotes the effective dimensionless pressure $\tilde{\mathcal{P}}$ at $t=10$ s.
@@ 4.2 Traveling wave solution
[…]
@@
$\psi(\tilde{\xi},\tilde{h})=e^{-2[\tilde{k}_{\rm s,{\rm i}}\tilde{\xi}+\tilde{K}_{z}(\tilde{h}-\tilde{h}^{\ast})]}$. (38)
@@
$\tilde{h}(\tilde{\rm w})=\tilde{h}^{\ast},\quad\tilde{J}(0)=\tilde{J}(\tilde{\rm w}),\quad\tilde{A}_{\rm d}=\int_{0}^{\tilde{\rm w}}\tilde{h}(\tilde{\xi})\,d\tilde{\xi}$. (41)
@@
We developed an iterative scheme to perform the numerical integration of Eq. (39), under the conditions given by Eq. (41). We start by guessing values of $\tilde{h}^{\prime\prime}(0)$, $\tilde{U}$ and $\tilde{\xi}_{\rm f}$ to perform the integration of Eq. (39) and modify them accordingly until these conditions are satisfied within a small relative error (typically, $10^{-7}$ or even smaller). After obtaining convergence for a given $\tilde{A}$, we use these converged values of $\tilde{h}^{\prime\prime}(0)$, $\tilde{U}$ and $\tilde{\xi}_{\rm f}$ as new guess values for $\tilde{A}+\Delta\tilde{A}$ (numerical continuation; $\Delta\tilde{A}$ has to be very small to reach convergence).
@@
This approximate formulation shows that the simple traveling-wave solution can capture, at least qualitatively, certain features of the experimental and numerical drops. Clearly, this description cannot account for the rear region of the drop (see Figs. 6 and 10(a)) because the values of $h^{\prime}$ and $h^{\prime\prime\prime}$ at the left contact line cannot be determined a priori, and both have been set to zero (for lack of better choice) in the traveling wave calculations.
Figure 10: (a) Thickness profile of the traveling wave solution for: $A=4.5$ nm, 66 nm, 88 nm, $10$ nm (b) Front speed, ${\rm v}_{\rm f}$ (mm/s, blue line), drop width, $\rm w$ ($m$, black line), and maximum height, $h_{\rm max}$ ($m$, red line), of the traveling drop as a function of AA; front speed is plotted on the left $y$-axis while drop width and maximum height are on the right $y$-axis.